OpenStudy (anonymous):

ill give medal What is the value of the variable that makes the statement true? a + (–7) + 3 = 0 A. –4 B. –3 C. 4 D. 7

OpenStudy (bishoppatton):

4

KjSaif (kjsaif):

what do you think it is i know the answer but want you to learn

KjSaif (kjsaif):

shotgun

OpenStudy (bishoppatton):

because 4 - (-7) = -3 -3 + 3 = 0

KjSaif (kjsaif):

well i guess he explained for you

KjSaif (kjsaif):

he is right

OpenStudy (bishoppatton):

And that is your answer

OpenStudy (anonymous):

sorry i needed to go to restroom

OpenStudy (bishoppatton):

ok